OpenStudy (anonymous):

Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

combine like terms

OpenStudy (uri):

-3x - 8 + 6x = 7 -3x+6x=7+8

OpenStudy (anonymous):

-3x-8+6x=7 +8 +8 -3x=15 now just divide and youll get x

OpenStudy (uri):

@ysenia Options?

OpenStudy (uri):

3x =15 x=5 ans.